Consider a simple first-order reaction AB. If you start with solution of 0.1 mol/L A, it takes 20 seconds to go to a concentration of 0.025 mol/L of A. (a) What is the value of the rate constant for this reaction? (b) Calculate half-life of the reaction. Select one or more: Rate constant = 0.0693 Rate constant = 0.6930 O Rate constant = 6.9310 Rate constant = 0.0250 Rate constant = 0.1250 O Rate constant has another value Half-life = 5 sec O Half-life = 10 sec Half-life = 20 sec Half-life = 40 sec Half-life = 50 sec O Half-life has another value


1
Expert's answer
2022-04-14T04:22:03-0400

a) Rate = dC/dt = (0.1 - 0.025) mol/L / 20 sec = 0.00375 mol/L s

b) 0.1/0.025 = 4 - 20 sec

0.1/0.05 = 2 - 10 sec

t1/2 = 10 sec